### v1

# Skew 2-Dyck paths via the kernel method

Abstract.
We continue on a recent concept introduced by Kariuki and Okoth, about skew 2-Dyck paths, introducing an additional down-step $L$, together with the usual steps $U$ (up) and $D$ down. There is the syntactical condition that ULUL and LULU can never occur. An automaton that checks these conditions is introduced, and the relevant generating functions are obtained by applying the kernel method to three functional equations. It is briefly discussed how the setting can be extended to $t$-Dyck paths. As a benefit, prefixes of skew $t$-Dyck paths are also enumerated.
Key words and phrases:
Dyck path, skew 2-Dyck paths, Generating functions, Kernel method
2020 Mathematics Subject Classification:
05A15
1. 2-Dyck paths
Dyck paths consist of up-steps $U=(1,1)$ and down-steps D=(1,1)D=(1,-1), start at the origin, never go below the $x$-axis and end at the $x$-axis. If a Dyck path does not necessarily end at the $x$-axis we call it a prefix of a Dyck path.
A standard reference on the enumeration of lattice paths is [1].
The 2-Dyck paths in the section title consist of up-steps $U=(1,1)$ and down-steps D=(1,2)D=(1,-2), never go below the $x$-axis and return to the $x$-axis. They will be the main subject of this paper, together with the epitheton skew. To clarify, we present a list of all 12 2-Dyck paths of 9 steps (length 9).
Figure 1. All 12 2-Dyck paths of 9 steps.
The new paper by Kariuki and Okoth [6] offers the new idea to introduce left down-steps. If L=(1,2)L=(-1,-2) this is not very interesting, as, related to the first paper on skew Dyck paths by Deutsch et al. [2] one should be careful that the path does not intersect (overlaps) itself, which in such a setting it wouldn’t do. Thus it is the right idea to stretch down-steps to D=(2,2)D=(2,-2) and L=(2,2)L=(-2,-2). Note that down-steps resp. down-left-steps still only count as one step. As we did before in our generating function based analysis of skew Dyck paths [3], instead of (2,2)(-2,-2) we draw (2,2)(2,-2) in red color. The syntactic conditions to be fulfilled are that ULUL and LULU are forbidden. A step $L$ can only follow a $D$-step, and cannot be followed by an $U$-step; note that the very last step may be an $L$-step.
Figure 2. All 12 2-Dyck paths of 9 steps, down-steps stretched.
Figure 3. All 2-Dyck paths of 9 steps, with possible red down steps indicated. This leads to altogether 19 paths.
The rest of this paper is devoted to an analysis (=enumeration) of skew 2-Dyck paths, using the kernel method. Some recent applications of these techniques can be found in [4], [5], [3]. There are some advantages: The analysis is in a sense straight forward; one does not need to invent clever decompositions of the paths in question, and one gets the enumeration of partial skew 2-Dyck paths as a bonus, as we have have all three generating functions leading to level $i$ by an up-step, down-step, red down-step in an explicit form.
In a last section, we indicate how to extend the concept from skew 2-Dyck paths to skew $t$-Dyck paths, again handling prefixes as a bonus.
2. Automaton to syntactically check 2-Dyck paths with left (red) steps
We start by constructing an automaton that checks whether the syntacting conditions are still fulfilled, i. e., ULUL and LULU are forbidden. According to the three types of steps, we have three layers of states (current level).
01122334455667788
Figure 4. Graph (automaton) to recognize skew 2-Dyck paths.
For the enumeration, we introduce $f_{i}=f_{i}(z)$, $g_{i}=g_{i}(z)$, $h_{i}=h_{i}(z)$, the generating functions of skew 2-Dyck paths leading to state $i$ in the first (second, third) layer, where the exponent $z^{k}$ counts the number of steps when starting at the origin (thick node on the first layer). The recursions are as follows:
$f_{0}$ $=1,\quad f_{i+1}=zf_{i}+zg_{i},\ i\geq 0,$
$g_{i}$ $=zf_{i+2}+zg_{i+2}+zh_{i+2},\ i\geq 0,$
$h_{i}$ $=zg_{i+2}+zh_{i+2},\ i\geq 0.$
They are easy to understand by distinguishing the last possible step that leads to level ii. Further, one introduces bivariate generating functions
$F(u)=F(u,z)=\sum_{i\geq 0}u^{i}f_{i},\quad G(u)=G(u,z)=\sum_{i\geq 0}u^{i}g_{i},\quad H(u)=H(u,z)=\sum_{i\geq 0}u^{i}h_{i}.$
By summing the recursions, we find
$F-1$ $=uz(F+G),$
$u^{2}G$ $=z(F-1-uf_{1})+z(G-g_{0}-ug_{1})+z(H-h_{0}-uh_{1}),$
$u^{2}H$ $=z(G-g_{0}-ug_{1})+z(H-h_{0}-uh_{1}).$
The solution is (obtained with the help of a computer)
$F$ $=\frac{z^{2}f_{1}u^{3}+z^{2}g_{1}u^{3}+z^{2}h_{1}u^{3}-u^{3}+z^{2}u^{2}+z^{2}g_{0}u^{2}+z^{2}h_{0}u^{2}-z^{3}f_{1}u+2zu-z^{3}}{zu^{4}-u^{3}-z^{2}u^{2}+2zu-z^{3}},$
[…]
First, we need to understand the denominator:
$zu^{4}-u^{3}-z^{2}u^{2}+2zu-z^{3}=z(u-s_{1})(u-s_{2})(u-s_{3})(u-s_{4}),$
with
$s_{1}$ $={\frac{1}{2}}{z}^{2}+{\frac{3}{16}}{z}^{5}+{\frac{17}{128}}{z}^{8}+{\frac{29}{256}}{z}^{11}+{\frac{861}{8192}}{z}^{14}+{\frac{6675}{65536}}{z}^{17}+{\frac{13231}{131072}}{z}^{20}+{\frac{52939}{524288}}{z}^{23}+\dots,$
$s_{2}$ $=\sqrt{2}\sqrt{z}+\frac{1}{4}{z}^{2}+{\frac{21}{64}}\sqrt{2}{z}^{7/2}+{\frac{29}{32}}{z}^{5}+{\frac{10727}{8192}}\sqrt{2}{z}^{13/2}+{\frac{1007}{256}}{z}^{8}+\dots,$
$s_{3}$ $=-\sqrt{2}\sqrt{z}+\frac{1}{4}{z}^{2}-{\frac{21}{64}}\sqrt{2}{z}^{7/2}+{\frac{29}{32}}{z}^{5}-{\frac{10727}{8192}}\sqrt{2}{z}^{13/2}+{\frac{1007}{256}}{z}^{8}$-$\dots$
$s_{4}$ $=\frac{1}{z}-{z}^{2}-2{z}^{5}-8{z}^{8}-39{z}^{11}-210{z}^{14}-1203{z}^{17}-7192{z}^{20}-44362{z}^{23}-280250{z}^{26}$-$\dots$
These 4 solutions where found using gfun and its algeqtoseries construction. [^1] Thanks to Cyril Banderier for valuable hints.
None of the factors $u-s_{1}$, $u-s_{2}$, $u-s_{3}$ can appear since we expect power series solutions, so they must cancel out. Hence, in $F$, $G$, $H$ the factor $(u-s_{1})(u-s_{2})(u-s_{3})$ can be divided out from both, numerator and denominator, [^2] Such a reasoning belongs to the realm of the kernel method. with the result
$F(u)$ $=\frac{z^{2}(f_{1}+g_{1}+h_{1})-1}{z(u-s_{4})},$
$G(u)$ $=\frac{-z^{2}(f_{1}+g_{1}+h_{1})}{z(u-s_{4})},$
$H(u)$ $=\frac{-z^{2}(g_{1}+h_{1})}{z(u-s_{4})}.$
Now one can plug in $u=0$, to find
$1=f_{0}$ $=\frac{z^{2}(f_{1}+g_{1}+h_{1})-1}{z(-s_{4})},$
$g_{0}$ $=\frac{-z^{2}(f_{1}+g_{1}+h_{1})}{z(-s_{4})},$
$h_{0}$ $=\frac{-z^{2}(g_{1}+h_{1})}{z(-s_{4})}.$
Simplifying the system leads to
$1+g_{0}=\frac{1}{zs_{4}},\quad g_{0}-h_{0}=\frac{z^{2}(1+g_{0})}{s_{4}}$
and further to
$g_{0}$ $=\frac{1}{zs_{4}}-1={z}^{3}+3{z}^{6}+13{z}^{9}+66{z}^{12}+365{z}^{15}+2131{z}^{18}+12921{z}^{21}+\dots,$
$h_{0}$ $=\frac{1}{zs_{4}}-\frac{z}{s_{4}^{2}}-1={z}^{6}+6{z}^{9}+34{z}^{12}+198{z}^{15}+1191{z}^{18}+7364{z}^{21}+\dots,$
$1+g_{0}+h_{0}=1+{z}^{3}+4{z}^{6}+\mathbf{19}{z}^{9}+100{z}^{12}+563{z}^{15}+3322{z}^{18}+20285{z}^{21}+\dots\;.$
The last series is the generating function of all skew 2-Dyck paths, and the number 19 refers to all such paths of length 9 as given in the examples.
Note that
$g_{1}+h_{1}=\frac{h_{0}s_{4}}{z},\quad f_{1}=z+zg_{0},$
which was used in the simplifications. Now, we can continue:
$F(u)=\frac{z^{2}f_{1}+z^{2}g_{1}+z^{2}h_{1}-1}{z(u-s_{4})}=\frac{z^{3}(1+g_{0})+zh_{0}s_{4}-1}{z(-s_{4})}\frac{1}{1-u/s_{4}},$
$G(u)=\frac{-z(f_{1}+g_{1}+h_{1})}{u-s_{4}}=\frac{-z^{2}(1+g_{0})-h_{0}s_{4}}{u-s_{4}}=\frac{z^{2}(1+g_{0})+h_{0}s_{4}}{s_{4}}\frac{1}{1-u/s_{4}},$
$H(u)=\frac{-z(g_{1}+h_{1})}{u-s_{4}}=\frac{-h_{0}s_{4}}{u-s_{4}}=h_{0}\frac{1}{1-u/s_{4}}.$
Consequently
$f_{k}$ $=[u^{k}]F(u)=\frac{-z^{3}(1+g_{0})-zh_{0}s_{4}+1}{z}s_{4}^{-k-1},$
$g_{k}$ $=[u^{k}]G(u)=(z^{2}(1+g_{0})+h_{0}s_{4})s_{4}^{-k-1},$
$h_{k}$ $=[u^{k}]H(u)=h_{0}s_{4}^{-k}.$
These are the promised generating functions of partial skew 2-Dyck paths, leading to level $k$ in all three possible ways.
The following comments might be useful. The denominator $zu^{4}-u^{3}-z^{2}u^{2}+2zu-z^{3}$ means that
$zf_{n}-f_{n+1}-z^{2}f_{n+2}+2zf_{n+3}-z^{3}f_{n+4}=0$
and the other sequences $g_{n}$, $h_{n}$ satisfy the same recursion.
Surprisingly, the numbers of skew 2-Dyck paths of length $3n$ have a simple generating function (seq A007564):
$R(z)=\frac{1}{6z}+\frac{1}{3}-\frac{\sqrt{1-8z+4z^{2}}}{6z}=1+z+4z^{2}+19z^{3}+100z^{4}+562z^{5}+3304z^{6}+20071z^{7}+\dots$
The extraction of coefficient is easy when using the substitution $z=\dfrac{t}{(1+t)(1+3t)}$; then $R=1+t$, and contour integration (or the Lagrange inversion formula) leads to
$[z^{n}]R(z)$ $=\frac{1}{2\pi i}\oint\frac{dz}{z^{n+1}}R$
$=\frac{1}{2\pi i}\oint dt\frac{1-3t^{2}}{(1+t)^{2}(1+3t)^{2}}\frac{(1+t)^{n+1}(1+3t)^{n+1}}{t^{n+1}}(1+t)$
$=\frac{1}{2\pi i}\oint\frac{dt}{t^{n+1}}[(1+t)-t(1+3t)]{(1+t)^{n}(1+3t)^{n-1}}$
$=\frac{1}{2\pi i}\oint\frac{dt}{t^{n+1}}(1+t){(1+t)^{n}(1+3t)^{n-1}}-\frac{1}{2\pi i}\oint\frac{dt}{t^{n+1}}t(1+3t){(1+t)^{n}(1+3t)^{n-1}}$
$=[t^{n}](1+t)^{n+1}(1+3t)^{n-1}-[t^{n-1}](1+t)^{n}(1+3t)^{n}$
$=\sum_{i=0}^{n}3^{i}\binom{n-1}{i}\binom{n+1}{n-i}-\sum_{i=0}^{n-1}3^{i}\binom{n}{i}\binom{n}{n-1-i}$
$=\sum_{i=0}^{n}3^{i}\binom{n-1}{i}\binom{n+1}{i+1}-\sum_{i=0}^{n-1}3^{i}\binom{n}{i}\binom{n}{i+1}$
$=\frac{1}{n}\sum_{i=0}^{n-1}3^{i}\binom{n}{i}\binom{n}{i+1}$.
The products of binomial coefficients may be written in terms of Narayana numbers $\frac{1}{n}\binom{n}{i}\binom{n}{i+1}$, compare [6].
The enumeration of 2-Dyck paths (without the skew-extension) leads to a third order algebraic equation, which, however, can be handled with the Lagrange inversion formula, see e. g., the thesis of Selkirk [7].
3. Generalization: skew $t$-ary Dyck-paths
We start with skew 3-Dyck paths, from which the general case about skew $t$-Dyck paths will become clear almost immediately. There is the usual up-step $U$, and a down-step $D$ of 3 units, and a red down-step $L$ of 3 units; the conditions are that ULUL and LULU are forbidden.
01122334455667788
Figure 5. Graph (automaton) to recognize skew 3-Dyck paths.
By setting up recursions and summing them, we get
$F-1$ $=uz(F+G),$
$u^{3}G$ $=z(F-1-uf_{1}-u^{2}f_{2})+z(G-g_{0}-ug_{1}-u^{2}g_{2})+z(H-h_{0}-uh_{1}-u^{2}h_{2}),$
$u^{3}H$ $=z(G-g_{0}-ug_{1}-u^{2}g_{2})+z(H-h_{0}-uh_{1}-u^{2}h_{2}).$
The solution is
$F$ $=\frac{u^{5}(z^{2}f_{2}+z^{2}g_{2}+z^{2}h_{2}-1)+\mathsf{something}}{zu^{6}-u^{5}-z^{2}u^{3}+2zu^{2}-z^{3}},$
$G$ $=\frac{-u^{5}z^{2}(f_{2}+g_{2}+h_{2})+\mathsf{something}}{zu^{6}-u^{5}-z^{2}u^{3}+2zu^{2}-z^{3}},$
$H$ $=\frac{-u^{5}z^{2}(g_{2}+h_{2})+\mathsf{something}}{zu^{6}-u^{5}-z^{2}u^{3}+2zu^{2}-z^{3}}.$
The denominator factors into 5 bad and one good root,
$zu^{6}-u^{5}-z^{2}u^{3}+2zu^{2}-z^{3}=z(u-s_{1})(u-s_{2})(u-s_{3})(u-s_{4})(u-s_{5})(u-s_{6}).$
The useful root is
$s_{6}=\frac{1}{z}-z^{3}-3z^{7}-16z^{11}-104z^{15}-749z^{19}-5748z^{23}-46069z^{27}-381093z^{31}-\dots\;.$
The factor $(u-s_{1})(u-s_{2})(u-s_{3})(u-s_{4})(u-s_{5})$ can be divided out from numerator resp. denominator, with the result
$F(u)$ $=\textsf{something}\;\frac{1}{1-u/s_{6}},$
$G(u)$ $=\textsf{something}\;\frac{1}{1-u/s_{6}},$
$H(u)$ $=\textsf{something}\;\frac{1}{1-u/s_{6}},$
and therefore
$f_{k}=[u^{k}]F(u)$ $=\textsf{something}\;s_{6}^{-k},$
$g_{k}=[u^{k}]G(u)$ $=\textsf{something}\;s_{6}^{-k},$
$h_{k}=[u^{k}]H(u)$ $=\textsf{something}\;s_{6}^{-k}.$
It is now clear how this translates to skew $t$-ary Dyck paths:
$f_{k}=[u^{k}]F(u)$ $=\textsf{something}\;s_{2t}^{-k},$
$g_{k}=[u^{k}]G(u)$ $=\textsf{something}\;s_{2t}^{-k},$
$h_{k}=[u^{k}]H(u)$ $=\textsf{something}\;s_{2t}^{-k};$
$s_{2t}$ is the only useful root of an algebraic equation of order $2t$ (in the variable $u$).
Perhaps some eager students might like to write programs so that the relevant results pop out when one enters the value of $t$.